OpenStudy (mitchi35):

Graph the circle with center (-1,2) and diameter 12. Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (mitchi35):

I don't understand the diameter 12. I know how to label. I just don't understand what it means to write the equation of the circle. Help please?

Vocaloid (vocaloid):

equation of a circle: (x-h)^2 + (y-k)^2 = radius^2 where (h,k) = (-1,2)

OpenStudy (mitchi35):

Thank you!

OpenStudy (mitchi35):

Oh wait. I still need help with finding the other four points to make the circle. Like how do I do that?

Vocaloid (vocaloid):

start plugging in 4 x-values, get 4-matching y values, then label those 4 points

OpenStudy (welshfella):

or to find the 4 points, since the radius of the circle = 12/2 = 6 you can add (6,0) to the center (-1,2) giving the point (5,2) also add (-6,0) giving (-7,2)

OpenStudy (welshfella):

and also ad (0 , 6) and (0 , -6)

OpenStudy (mitchi35):

Okay.Thanks guys/girls! :)